OpenStudy (anonymous):

find the dervative of y=(lnx+e^x)^5

OpenStudy (helder_edwin):

you have to use the chain rule do u know this

OpenStudy (anonymous):

Not really

OpenStudy (lgbasallote):

power rule + chain rule to be exact

OpenStudy (helder_edwin):

the chain rule says you first differentiate the function that is outside and then you inside little by little \[ \large (f\circ g)'=(f'\circ g)\cdot g' \]

OpenStudy (helder_edwin):

here you have \[ \large f(x)=\ln x \] \[ \large g(x)=e^x \] \[ \large h(x)=x^5 \] then \[ \large y(x)=(h\circ(f+g))(x) \]

OpenStudy (helder_edwin):

do you know how to find h'(x) ?